OpenStudy (anonymous):

y = 1 ( x + 3 )^2 - 25

OpenStudy (anonymous):

So far i've got y=1x^2-16

OpenStudy (anonymous):

1x+3^2-25

geerky42 (geerky42):

None.

OpenStudy (anonymous):

1x+9-25

OpenStudy (anonymous):

why wouldn't the ^2 apply to 1x?

OpenStudy (anonymous):

3x3=9

geerky42 (geerky42):

\((x+3)^2 = (x+3)(x+3) = x^2 + 3x + 3x + 3^2 = x^2+6x+9\)

OpenStudy (anonymous):

u got rid of that wen multiply by 1

OpenStudy (anonymous):

3x1=3

geerky42 (geerky42):

Do you ever know what you are doing? @kittycat123

OpenStudy (anonymous):

1timesx 1x or x

OpenStudy (anonymous):

yes

geerky42 (geerky42):

Question is "y = 1 ( x + 3 )^2 - 25" Note "^2"

OpenStudy (anonymous):

y=(x+8)(x-2)

OpenStudy (anonymous):

where did you get 8..?

geerky42 (geerky42):

When you simplify completely, then you factor it, you should get \((x+8)(x-2)\)

OpenStudy (anonymous):

okay, thank you for all of your help !:)